# LaB-RAG: Label Boosted Retrieval Augmented Generation for Radiology Report Generation

### v1

@@ 5.3 Extended Analyses
Motivated by the results comparing LaB-RAG to literature (Fig. 3), we sought to analyze why we were comparable to SOTA in F1RadGraph for “Impression” generation but not “Findings” generation. Tab. 2 includes an example of an actual “Findings” section and its corresponding generations by LaB-RAG and CXRMate [57]. We annotate the entities identified by CheXbert [70] with their label. These entities are also identified by RadGraph [27], which also identifies many more. Examining RadGraph annotations, we observe that the entities of the actual report more closely align with those of CXRMate’s generation, hence the F1RadGraph achieved for CXRMate was $0.6$ vs Lab-RAG’s $0.2$. However, we also observe that the computed CheXbert labels of both LaB-RAG and CXRMate exactly match those of the actual report, resulting in F1CheXbert of $1.0$.
We argue that F1CheXbert overvalues these generated “Findings” because it relies solely on 14 labels that describe disease related findings. This is most relevant to novel composite RRG metrics, such as RadCliQ [79], which include CheXbert as a component metric. As LaB-RAG explicitly uses these same labels in its label filter and format, we effectively ensure that these labels will be present in the generated report. This may also be the reason for LaB-RAG’s improved F1RadGraph performance in “Impression” generation. Since the “Impression” describes summary diagnoses, CheXbert and RadGraph entities may be more closely aligned with each other. It is thus an open question of which label set to use for label boosting. An interesting possibility may be using RadGraph derived labels.
5.4 Other Limitations
Though we make no claim of LaB-RAG’s clinical viability, an analysis of hallucinations is an important step towards this discussion for any model; Tab. 2 certainly shows hallucinations in both generations. Another potential limitation arises from the CheXpert [25] labels included by MIMIC-CXR-JPG [32]; these labels were derived on mixed report sections, however we do not consider their source for our section-specific generation. Lastly, a limitation of all methods using pretrained LLMs is the difficulty of detecting data leakage from pretraining over unknown training data.
@@ 6 Conclusion
In this work, we present LaB-RAG: a modular framework for image captioning using pretrained models. We study and analyze LaB-RAG in the context of RRG, showing that it achieves SOTA on clinical language metrics. We offer insights into the importance of different components of LaB-RAG, while also critiquing a widely used RRG metric, arguing we can inflate its results. Throughout our experiments, we stress the simplicity of LaB-RAG as an ensemble of frozen DL models boosted by classical linear models. The key to LaB-RAG is leveraging categorical labels as natural descriptors of images. Doing so enables LaB-RAG to use off-the-shelf models which do not need to be jointly trained, a task which requires large paired datasets. This enables LaB-RAG to be fully modular and flexible for users.
Supplementary Material
S1 Abbreviations and Terms
@@
Table S1: Reference of abbreviations.
S2 Dataset Extended Information
@@
MIMIC-CXR [35] imaging studies were collected in the emergency department at the Beth Israel Deaconess Medical Center (BIDMC) in Boston, MA between 2011 and 2016. Multiple chest X-rays may be present for a single study, and a patient may have multiple imaging studies. The training, validation, and testing splits provided by MIMIC-CXR are split at the patient level, meaning all images for all studies belonging to a single patient are contained in one split. The demographic breakdown of studies in each split is presented in Tab. S2, though the final number of patients and imaging studies considered in each of our experiments depends on the availability of all required data modalities. Demographic information was collected from the MIMIC-IV v2.2 [37, 31] hosp module and MIMIC-IV-ED v2.2 [30]. All data were accessed through PhysioNet [20].
For images, we select one image per study by preferentially selecting images based on the captured view position. The number of selected image views per split is listed in Tab. S3. We follow the order of preference for view positions from the ACL 2024 BioNLP workshop’s RRG shared task [15]; the exact list can be found in Tab. S4. We use code inspired by the MIMIC Code Repository [34] to extract radiology report sections from the free text reports. The number of extracted findings and impression sections per split is listed in Tab. S5.
Ethical Considerations for Data Usage. As the MIMIC [37] datasets were collected from the BIDMC, the BIDMC institutional review board (IRB) granted a waiver of informed consent and approved the collection and sharing of the data for research purposes following proper de-identification according to HIPAA. As users of the data, we follow procedures outlined by PhysioNet [20] to have all researchers with data access complete human research training and sign a data use agreement (DUA). We strictly control access to the data in our compute environments and make no efforts to re-identify data subjects from any free-text data (which may have escaped de-identification) or any other sources. Additionally, as we do no fine-tuning of our own LLMs, we do not need to consider whether it is a DUA violation to share LLM weights which may have memorized and can reproduce training text.
S3 LaB-Classifier Training Details
@@
We fit 14 per-label binary logistic regressions over the unprojected 512d image embeddings from BioViL-T (see Sec. 3.2). We derive ground-truth labels to train our models with by binarizing the CheXpert labels as positive (value of 1) or not (value of 0, -1, or null). We fit the models on the training split and find a per-label probability threshold which maximizes the f1-score over the validation split. The receiver operator characteristic and precision recall curves of the 14 labels over the test split are presented in Fig. S2. We adopt the same heuristic as above for computing the “Other” label for the image. For training the logistic classifiers, we use sklearn [61] v1.5.1 with L2 regularization, the saga solver for 500 iterations, and otherwise default hyperparameters.
Figure S1: Overview of training per-label LaB-Classifier logistic regressions.
Figure S2: Test set ROC and PR curves of LaB-Classifier, a set of 14 per-label logistic regressions.
S4 Evaluation Strategy Extended Details
@@
Overall Train Val Test
Count, N Patients 65379 64586 500 293
Studies 227835 222758 1808 3269
Age Median [Q1,Q3] 65 [52,76] 65 [52,76] 63 [52,77] 69 [61,77]
Missing, N (%) 11825 (5.2) 11522 (5.2) 134 (7.4) 169 (5.2)
Sex, N (%) Female 105117 (48.7) 102951 (48.7) 737 (44.0) 1429 (46.1)
Male 110893 (51.3) 108285 (51.3) 937 (56.0) 1671 (53.9)
Missing 11825 (5.2) 11522 (5.2) 134 (7.4) 169 (5.2)
Race, N (%) White 141531 (65.5) 138343 (65.5) 1081 (64.6) 2107 (68.0)
Black 35788 (16.6) 34920 (16.5) 186 (11.1) 682 (22.0)
Hispanic or Latino 12358 (5.7) 12150 (5.8) 110 (6.6) 98 (3.2)
Other 9088 (4.2) 8857 (4.2) 190 (11.4) 41 (1.3)
Unknown 8430 (3.9) 8343 (4.0) 51 (3.0) 36 (1.2)
Asian 7927 (3.7) 7753 (3.7) 47 (2.8) 127 (4.1)
AIAN 644 (0.3) 628 (0.3) 7 (0.4) 9 (0.3)
NHPI 205 (0.1) 203 (0.1) 2 (0.1) 0 (0.0)
Missing 11864 (5.2) 11561 (5.2) 134 (7.4) 169 (5.2)
Table S2: Per-study demographic breakdown of MIMIC defined training/validation/testing splits. AIAN = American Indian or Alaska Native, NHPI = Native Hawaiian or Pacific Islander.
View, N (%) Overall Train Val Test
PA 85867 (37.7) 84323 (37.9) 661 (36.6) 883 (27.0)
AP 132264 (58.1) 129034 (57.9) 1072 (59.3) 2158 (66.0)
LATERAL 176 (0.1) 172 (0.1) 2 (0.1) 2 (0.1)
LL 2225 (1.0) 2160 (1.0) 17 (0.9) 48 (1.5)
AP AXIAL 1 (0.0) 1 (0.0) 0 (0.0) 0 (0.0)
LAO 1 (0.0) 1 (0.0) 0 (0.0) 0 (0.0)
LPO 1 (0.0) 1 (0.0) 0 (0.0) 0 (0.0)
Table S3: Count of selected single view per-study. PA = postero-anterior, AP = antero-posterior, LL = left lateral, LAO = left anterior oblique, LPO = left posterior oblique.
Rank View Rank View Rank View
1 PA 6 AP LLD 11 RAO
2 AP 7 AP RLD 12 LPO
3 LATERAL 8 PA RLD 13 XTABLE LATERAL
4 LL 9 PA LLD 14 SWIMMERS
5 AP AXIAL 10 LAO 15 (Empty String)
Table S4: Order of view preference for selecting single image for multi-image studies. PA = postero-anterior, AP = antero-posterior, LL = left lateral, LLD = left lateral decubitus, RLD = right lateral decubitus, LAO = left anterior oblique, RAO = , LPO = left posterior oblique.
Section, N (%) Overall Train Val Test
Impression 189561 (83.2) 185816 (83.4) 1521 (84.1) 2224 (68.0)
Findings 155716 (68.3) 152173 (68.3) 1196 (66.2) 2347 (71.8)
Both 128032 (56.2) 125417 (56.3) 991 (54.8) 1624 (49.7)
Table S5: Count of reports with extracted “Findings”, “Impression”, or both sections.
S5 Prompts
Prompt Type Prompt Text
Naive You are an expert radiological assistant. Your task is to generate a radiology report after <<Report>> given context information. The context information contains examples of reports written for similar cases. Use the examples to generate a report for the current case. Strictly follow the instructions below to generate the reports. *Instructions** 1. The report must be based on the information in the context. 2. The report must mimic the style of the reports shown in the context. 3. Do not generate blank reports. CONTEXT: Example: 1 (Report Text) (More Examples) Now generate the report for the current case. Always generate reports based on the examples shown. <<Report>>
Table S6: The “Naive” prompt does not incorporate labels.
Prompt Type Prompt Text
Simple You are an expert radiological assistant. Your task is to generate a radiology report after <<Report>> given context information. The context information contains examples of reports written for similar cases and their associated labels. Use the examples and their associated labels to generate a report for the current case based on the current label. Strictly follow the instructions below to generate the reports. *Instructions** 1. The report must be based on the information in the context and the current label. 2. The report must mimic the style of the reports shown in the context. 3. Do not generate blank reports. CONTEXT: Example: 1 Label: (Positive Labels) (Report Text) (More Examples) Now generate the report for the current case using its label below. Always generate reports based on the examples shown. Label: (Positive Labels) <<Report>>
Table S7: The “Simple” prompt formats positive labels for the examples and target image.
Prompt Type Prompt Text
Verbose You are an expert radiological assistant. Your task is to generate a radiology report after <<Report>> given context information. The context information contains examples of reports written for similar cases and their associated labels. The labels provided are expert annotations. More information about the labels is described below. The individual labels used represent common chest radiographic observations and fall under four categories: ‘Positive’, ‘Negative’, ‘Uncertain’ and ‘Unmentioned’. These categories correspond to the mention or presence of the labels or their equivalent in the report. Below is a description and example of each of the label categories: 1. ‘Positive’: A label is positive if the associated observation or disease is stated as present        in the report, for example: ‘moderate bilateral effusions observed’. 2. ‘Negative’: A label is negative if the associated observation or disease is stated as absent        in the report, for example: ‘no evidence of pulmonary edema’. 3. ‘Uncertain’: A label is uncertain if there is ambiguity about the presence or absence of        the associated observation or disease in the report, for example: ‘pneumonia cannot be excluded        in the appropriate clinical context’. 4. ‘Unmentioned’: A label is unmentioned if there is no mention of the associated observation        or disease in report. Use the examples, their associated labels, and the label descriptions to generate a report for the current case based on the current label. Strictly follow the instructions below to generate the reports. *Instructions** 1. The report must be based on the information in the context and the current label. 2. The report must mimic the style of the reports shown in the context. 3. Do not generate blank reports. CONTEXT: Example: 1 Positive: (Positive Labels) Negative: (Negative Labels) Uncertain: (Uncertain Labels) Unmentioned: (Unmentioned Labels) (Report Text) (More Examples) Now generate the report for the current case using its label below. Always generate reports based on the examples shown. Positive: (Positive Labels) Negative: (Negative Labels) Uncertain: (Uncertain Labels) Unmentioned: (Unmentioned Labels) <<Report>>
Table S8: The “Verbose” prompt formats positive, negative, uncertain, and unmentioned labels for the examples and target image.
Prompt Type Prompt Text
Instruct (Same as Verbose) *Instructions** 1. The report must be based on the information in the context and the current label. 2. The report must mimic the style of the reports shown in the context. 3. Do not generate blank reports. 4. Ensure that the positive labels are clearly described as being present in the report,        using example language from the context. 5. Ensure that the negative labels are clearly described as being absent in the report,        using example language from the context. 6. Describe the uncertain labels as necessary. 7. Ensure that the unmentioned labels are not mentioned in the report. (Same as Verbose)
Table S9: The “Instruct” prompt uses the same label format as “Verbose” and adds additional instructions.
S6 Literature Model Inference
@@ 5 AP AXIAL 10 LAO 15 (Empty String)
In this section, we provide our extended descriptions for baseline models from the literature. We select models based on model architecture, generated report section, SOTA performance on evaluation metrics, and finally availability of open source inference code. The final set of models we compare against are as follows: CXR-RePaiR [18], CXR-ReDonE [64], X-REM [28], RGRG [73], CheXagent [12], and CXRMate-RRG24 [57]. We compare models over the intersection of each method’s test split subsets.
@@
CXR-RePaiR formulates report generation as a pure retrieval task. It uses a model, trained via CLIP on radiology report-image pairs from MIMIC-CXR, to rank similarity between test images and a large retrieval corpus. We use the mode of CXR-RePaiR where the generated report output is the the top-1 retrieved whole reports.
@@
CXR-ReDonE improves upon CXR-RePaiR by preprocessing the training and retrieval reports to remove references to prior imaging studies. Additionally, the joint-embedding model was trained via ALBEF instead of CLIP. We use the provided checkpoint of the retrieval model which was trained over the new data and via the new method; we do inference using whole reports with priors omitted as the retrieval set. The generated reports are then the top-1 retrieved whole report. CXR-ReDonE uses the “Impression” sections preprocessed by CXR-RePaiR and so results in the same test split subset.
X-REM also trains a vision-language model via ALBEF, however they introduce a novel similarity metric during training which incorporates CheXbert labels. An intermediary step retrieves the top-10 whole reports ranked by their new similarity score. Finally, they apply a model-based natural language filter to each retrieved report and only select those that are deemed relevant up to a limit; we adopt the author default limit of 2 reports, thus the output report is a concatenation of up to 2 reports. X-REM also uses the “Impression” sections preprocessed by CXR-RePaiR with the same test split subset.
CheXagent is presented by its authors as a foundation model trained to follow instructions in the CXR domain. CheXagent follows a complex training scheme and utilizes other CXR datasets. At a high-level, CheXagent is trained by aligning image latent embeddings to the an LLM’s token embedding space. Additionally, the authors introduce a novel dataset for instruction tuning, their final training step.
@@
CXRMate uses an encoder-decoder transformer architecture. It is both able to generate a single report over multiple images from a single imaging study and it takes as input reports from prior studies. Additionally, it is trained with a complex reinforcement learning framework. Interestingly, the authors experiment with RadGraph and CXR-BERT in their reward function, arguing that CXR-BERT better captures radiology report semantics; CXR-BERT is the language encoder of the vision-langauge model BioViL [4].
@@
We specifically use the checkpoint of CXRMate submitted to the ACL 2024 BioNLP workshop’s task for RRG [57], though we only input the single image per study defined by our data preprocessing steps. While CheXagent generates both “Findings” and “Impression” jointly, the model provides a utility to split these sections after generation. We are then able to run inference using our subsets of the test split for which we have “Findings”, “Impression”, or both sections.

### v2

@@ 5.3 Extended Analyses
We sought to understand the difference in Figure 3 Left between SOTA measured by F1CheXbert and underperformance on F1RadGraph. Table 2 presents a real “Findings” section for a MIMIC-CXR report written by a radiologist and its corresponding generations by LaB-RAG and CXRMate (Nicolson et al., 2024b). We annotate entities which result in a specific CheXbert (Smit et al., 2020) label, namely “Cardiomegaly” and “Support Device”. As the computed CheXbert labels of both generations from LaB-RAG and CXRMate exactly match those of the actual report, this results in F1CheXbert of $1.0$. The CheXbert entities are also identified by RadGraph (Jain et al., 2021), however RadGraph identifies many more such entities. Examining RadGraph annotations, we observe that the entities of the actual report more closely align with those of CXRMate’s generation, hence the F1RadGraph achieved for CXRMate was $0.63$ vs Lab-RAG’s $0.48$. Yet, overall, as in Yu et al. (2023), both generated reports make substantive errors which may impact clinical interpretation.
@@ 6 Conclusion
We present LaB-RAG: a new modular framework for image captioning that leverages categorical labels extracted by small models to boost large language models. We study and analyze LaB-RAG in the context of RRG, showing that it achieves SOTA on clinical language metrics. We offer insights into the importance of different components of LaB-RAG, suggesting potential for future synergy with SFT and other training methods. The key to LaB-RAG is leveraging inexpensive models to derive categorical labels as natural descriptors of images. Doing so enables LaB-RAG to further boost models with strong capabilities within a flexible and modular framework.
Ethics Statement
For our study, though all data we use in our study are publicly available and deidentified, we do still work with real patient data. These data were retrospectively collected with no impact to real patient care. The dataset authors follow standardized procedures to de-identify data and remove protected health information (PHI) in accordance with HIPAA regulations. For MIMIC-CXR, we follow procedures outlined by PhysioNet (Goldberger et al., 2000) to have all researchers with data access complete human research training and sign the PhysioNet data use agreement (DUA). We strictly control access to the data in our compute environments in accordance with the credentialing process for data access. Though these data access steps are not required for CheXpert Plus (Chambon et al., 2024), we still follow this strict procedure to ensure best practices for data governance. We make no efforts to re-identify data subjects from any free-text data (which may have escaped de-identification) or any other sources. Additionally, as we do no fine-tuning of our own LLMs over this data, we do not need to consider whether it is a DUA violation to share LLM weights which may have memorized and can reproduce training text. To the best of our ability, we adhere to scientific principles of research integrity. We make no claims to the real-world clinical viability of any of our models and recognize that strict validation must be done before any such consideration can be made. These are necessary steps to protect the patients whom we aim to help.
Reproducibility Statement
Our code and instructions for full reproducibility is available at: https://github.com/uc-cdis/label-boosted-RAG-for-RRG.
Appendix
Appendix A Abbreviations and Terms
Table 3: Reference of abbreviations.
@@ Appendix A Abbreviations and Terms
Appendix B Extended Methods
B.1 Chest X-ray Datasets
We utilize two CXR datasets for our study, MIMIC-CXR (Johnson et al., 2019) and CheXpert Plus (Chambon et al., 2024). The descriptive statistics of the patient cohorts used in our studies is presented in Table 9. In our experiments, we use the training and validation splits (described below) as our retrieval set and evaluate inference results over the test split.
@@ B.1 Chest X-ray Datasets
MIMIC-CXR: Imaging studies were collected in the emergency department at the Beth Israel Deaconess Medical Center (BIDMC) in Boston, MA between 2011 and 2016. Multiple chest X-rays may be present for a single imaging study, and a patient may have multiple imaging studies. We use the training, validation, and testing splits provided by Johnson et al. (2019). The dataset is split at the patient level, meaning all images for all studies belonging to a single patient are contained in one split. Demographic information was joined from the MIMIC-IV v2.2 (Johnson et al., 2023b; a) hosp module. We use code from the MIMIC Code Repository (Johnson et al., 2018) to extract radiology report sections from the free text reports. All data were accessed through PhysioNet (Goldberger et al., 2000).
CheXpert Plus: Imaging studies were collected from Stanford Hospital in Palo Alto, CA between 2002 and 2017. CheXpert Plus is an enhancement of the original CheXpert dataset (Irvin et al., 2019) to include, among other new facets, radiology reports. As such, besides small amounts of missing data, the provided patient-level data splits between the two versions have remained largely unchanged. This is significant for our experimental design, as the dataset authors only define development and validation splits. Following Huang et al. (2021), we use the provided validation split as “test” and resample the development set into new “train” and “validate” splits. This better enables us to do multi-step experiments (see Section B.2) and prevent data-leakage. We do note that the test split of CheXpert Plus is both absolutely and relatively small with N=200 studies; the small test set size is further exacerbated when considering only studies with specific report sections (i.e. N=61 for CheXpert Plus test-set studies with “Findings” sections).
Multi-view studies: It is standard clinical practice to capture multiple views for a single imaging study, conceptually showing alternate angles of the patient. For our experiments, we select one image per study by preferentially selecting images based on the captured view position. Broadly, we select frontal views, then lateral views, then all other views. Specifically, we use the order of preference for view positions from the ACL 2024 BioNLP workshop’s RRG shared task (Demner-Fushman et al., 2024). The number of selected views per dataset and split is given in Table 8.
Extracting Labels: By default, we use CheXbert (Smit et al., 2020) extracted labels for experiments. For our study on label quality, we additionally use CheXpert (Irvin et al., 2019) extracted labels. We specifically extract labels from the ground-truth radiology report sections we aim to generate, i.e. Findings or Impression; we do not use the labels provided by either dataset authors as these were derived over mixed report sections. The CheXbert extracted label prevalence across datasets and splits is presented in Table 10.
B.2 LaB-Classifier Training
@@
We fit 14 per-label binary logistic regressions over extracted image embeddings (see Section 3.2). We derive “ground-truth” labels to train our classifiers by binarizing the extracted labels as positive (value of 1) or not (value of 0, -1, or null). We fit the models on the training split and find a per-label probability threshold which maximizes the f1-score over the validation split. We repeat this process for each alternate image embedding or labeler we experiment with. We present the per-label F1 score of the predicted labels across embeddings, labelers, and datasets in Table 11. We adopt the same heuristic as in Section 3.1 for computing the predicted image label, where if no positive label is predicted by the classifier, the image is assigned an implicit positive “Other” label. For training the logistic classifiers, we use sklearn (Pedregosa et al., 2011) v1.5.1 with L2 regularization, the saga solver for 500 iterations, and otherwise default hyperparameters. From a computational cost perspective, training and inference of 14 linear models is orders of magnitude fewer in parameters than a DL-based image encoder model.
Figure 5: Overview of training per-label LaB-Classifier logistic regressions.
B.3 Other Implementation Details
Image Embeddings: We extract zero-shot image embeddings with the frozen image encoders of each model. For BioViL-T and GLoRIA, we use projected embeddings for image-based retrieval (128d and 768d, respectively) and we use unprojected embeddings for training our image classifier (512d and 2048d, respectively). For ResNet50, we use the final 2048d hidden layer output for both retrieval and classifier training.
LLM Inference: We serve the generative language model component of LaB-RAG using vLLM (Kwon et al., 2023) and do greedy decoding up to 512 tokens, with a set seed, and temperature 0.
B.4 Evaluation Strategy
@@
B.5 Literature Model Inference
@@
In this section, we provide our extended descriptions for baseline models from the literature evaluated over MIMIC-CXR. We select models based on model architecture, generated report section, SOTA performance on evaluation metrics, and finally availability of open source inference code. The final set of models we compare against are as follows: CXR-RePaiR (Endo et al., 2021), CXR-ReDonE (Ramesh et al., 2022b), X-REM (Jeong et al., 2024), RGRG (Tanida et al., 2023), CheXagent (Chen et al., 2024b), and CXRMate-RRG24 (Nicolson et al., 2024b). We compare models over the intersection of each method’s test split subsets.
@@
CXR-RePaiR: formulates report generation as a pure retrieval task. It uses a model, trained via CLIP on radiology report-image pairs from MIMIC-CXR, to rank similarity between test images and a large retrieval corpus. We use the mode of CXR-RePaiR where the generated report output is the the top-1 retrieved whole reports.
@@
CXR-ReDonE: improves upon CXR-RePaiR by preprocessing the training and retrieval reports to remove references to prior imaging studies. Additionally, the joint-embedding model was trained via ALBEF instead of CLIP. We use the provided checkpoint of the retrieval model which was trained over the new data and via the new method; we do inference using whole reports with priors omitted as the retrieval set. The generated reports are then the top-1 retrieved whole report. CXR-ReDonE uses the “Impression” sections preprocessed by CXR-RePaiR and so results in the same test split subset.
X-REM: also trains a vision-language model via ALBEF, however they introduce a novel similarity metric during training which incorporates CheXbert labels. An intermediary step retrieves the top-10 whole reports ranked by their new similarity score. Finally, they apply a model-based natural language filter to each retrieved report and only select those that are deemed relevant up to a limit; we adopt the author default limit of 2 reports, thus the output report is a concatenation of up to 2 reports. X-REM also uses the “Impression” sections preprocessed by CXR-RePaiR with the same test split subset.
CheXagent: is presented by its authors as a foundation model trained to follow instructions in the CXR domain. CheXagent follows a complex training scheme and utilizes other CXR datasets. At a high-level, CheXagent is trained by aligning image latent embeddings to the an LLM’s token embedding space. Additionally, the authors introduce a novel dataset for instruction tuning, their final training step.
@@
CXRMate: uses an encoder-decoder transformer architecture. It is both able to generate a single report over multiple images from a single imaging study and it takes as input reports from prior studies. Additionally, it is trained with a complex reinforcement learning framework. Interestingly, the authors experiment with RadGraph and CXR-BERT in their reward function, arguing that CXR-BERT better captures radiology report semantics; CXR-BERT is the language encoder of the vision-langauge model BioViL (Boecking et al., 2022).
@@
We specifically use the checkpoint of CXRMate submitted to the ACL 2024 BioNLP workshop’s task for RRG (Nicolson et al., 2024b), though we only input the single image per study defined by our data preprocessing steps. While CheXagent generates both “Findings” and “Impression” jointly, the model provides a utility to split these sections after generation. We are then able to run inference using our subsets of the test split for which we have “Findings”, “Impression”, or both sections.